OpenStudy (kise_ryouta):

Solve for x. 2/3x=10

OpenStudy (iluvsoccer):

2/3x basically means 2/3 * x so now in order to get x on one side, we have to do the opposite. 2/3 / 10/1

OpenStudy (iluvsoccer):

2/3 * 1/10

OpenStudy (rose4825):

\[\frac{ 2 }{ 3 }x=10\] Get x on it's own by multiplying both sides by 3/2. \[\left( \frac{ 3 }{ 2 } \right)\frac{ 2 }{ 3 } x=\left( \frac{ 3 }{ 2 } \right)*10\] Multiply 3/2 and 10. To make this easier, make 10 a fraction by writing 10/1 \[x=\frac{ 3 }{ 2 }*\frac{ 10 }{ 1 }\] Multiply. \[x = \frac{ 30 }{ 2 }\] Simplify. \[x=15\]

OpenStudy (rose4825):

Does that make sense?

OpenStudy (kise_ryouta):

yes it does thanks and can you help me with more pls

OpenStudy (rose4825):

Sure, I'll see what I can do.

OpenStudy (kise_ryouta):

thanks ^^

OpenStudy (rose4825):

no problem :)

OpenStudy (kise_ryouta):

Two trains leave the same station at the same time, but are headed in opposite directions. One train travels at 70 mph and the other train travels at 80 mph. How much times passes until the trains are 600 mi apart?

OpenStudy (rose4825):

So basically the word problem contains the following equation: \[70x*80x=600\] Add. \[150x=600\] Get x on its own by dividing both sides by 150. \[(150)\div150x=\frac{ 600 }{ 150 }\] Simplify. \[x=4\] I don't know what the unit of time that the question is asking, but I would say hours based on the fact that the trains are traveling "mph" so that answer (in my opinion) would be 4 hours.

OpenStudy (rose4825):

But I am not too confident with this answer

OpenStudy (kise_ryouta):

hmm i can kinda see where your coming from

OpenStudy (kise_ryouta):

Solve for c. a=11+4b−4c @rose4825